# MambaLRP: Explaining Selective State Space Sequence Models

### v1

@@
Abstract
@@
  • Keywords:  Explainable AI, State Space Models, Mamba, Long-Range Dependencies
@@ 1 Introduction
Explaining the predictions of Mamba models is however challenging due to their highly non-linear and recurrent structure. A recent study [4] suggests viewing these models as attention-based models, enabling the use of attention-based explanation methods [1, 17]. Yet, the explanations produced by attention-based techniques are often unreliable and exposed to potential misalignment between input features and attention scores [65, 34]. As an alternative, Layer-wise Relevance Propagation (LRP) [9] decomposes the model function with the goal of explicitly identifying the relevance of input features by applying purposely designed propagation rules at each layer. A distinguishing feature of LRP is its adherence to a conservation axiom, which prevents the artificial amplification or suppression of feature relevance in the backward pass. LRP has been demonstrated to produce faithful explanations across various domains (e.g. [6, 3, 18]). Nevertheless, the peculiarities of the Mamba architecture are not addressed by the existing LRP procedures, which may lead to the violation of the conservation property and result in unreliable explanations.
@@ 2 Related Work
Transformers [59] have emerged as the most widely used architectures for sequence modeling. However, their computational limitations, particularly with large sequence lengths, have restricted their applicability in modeling long sequences. Addressing these computational limitations, recent works [30, 31] have introduced structured state space sequence models (SSMs) as an alternative approach. SSMs are a class of sequence modeling methods, leveraging the strengths of recurrent, convolutional, and continuous-time methods, demonstrating promising performance across various domains, including language [26, 40], image [67, 12, 44], and video [61] processing, and beyond [50, 19, 38]. A recent advancement by Gu and Dao [29] introduced selective SSM, an enhanced data-dependent SSM with a selection mechanism that adjusts its parameters based on the input. Built on this dynamic selection, the Mamba architecture fuses the SSM components with multilayer perceptron (MLP) blocks. This fusion simplifies the architecture while improving its ability to handle various sequence modeling tasks, including applications in language processing [5, 45, 62], computer vision [37, 73, 68], medical imaging [39, 66, 27, 47, 36, 64, 63] and graphs [60, 13]. This fast adoption of SSMs and Mamba models underscores the need for reliable explanations of their predictions.
@@
In efforts to explain Mamba models, Ali et al. [4] recently proposed viewing their internal computations as an attention mechanism. This approach builds upon previous works that use attention signal as explanation, including Attention Rollout [1] and variants thereof [17, 16]. While these approaches can provide some insight, they inherit the limitations of using attention as an explanation [65, 34], including their inability to capture potential misalignment between tokens and attention scores, and the limited performance in empirical faithfulness evaluations. Alternative Explainable AI methods, not yet applied to Mamba models but in principle applicable to any model, include techniques using input perturbations [71, 74, 25] or leveraging gradient information [10, 54, 58, 55, 53]. Despite their generality, these models have certain drawbacks, such as requiring multiple function evaluations for a single explanation or being susceptible to gradient noise, resulting in subpar performance, as our benchmark experiment will demonstrate. To tackle these challenges, we introduce MambaLRP as an efficient solution for the computation of reliable and faithful explanations that are theoretically grounded in the axiom of relevance conservation.
@@ 4.2 Relevance propagation in selective SSMs (S6)
Figure 2: Unfolded view of SSM, highlighting two subsets of nodes, the relevance of which should be conserved throughout relevance propagation.
@@ 4.3 Relevance propagation in multiplicative gates
As for the previous examples, this ensures the conservation property $\mathcal{R}(x)=\mathcal{R}(y)$. An alternative would have been to make $y$ linear by detaching only one of the terms in the product, as done for the SiLU activation or the SSM component. However, the strategy of Eq. (9) better maintains the directionality given by the gradient. We further compare these alternatives in an ablation study presented in Appendix C.5, demonstrating empirically that our proposed approach performs better.
@@ 4.4 Additional modifications and summary
The propagation strategies developed for the Mamba-specific components complement previously proposed approaches for other layers, including propagation through RMSNorm layers [3] and convolution layers via robust LRP-$\gamma$ rules [43, 22] and their generalized variants. A summary of these additional enhancements is provided in Appendix C.2.
Overall, our MambaLRP procedure can be implemented as a sequence of two steps:
@@ 5 Experiments
To evaluate our proposed approach, we benchmark its effectiveness against various methods previously proposed in the literature for interpreting neural networks. We empirically evaluate our proposed methodology using Mamba-130M and Mamba-1.4B language models [29], which are trained on diverse text datasets. For the vision experiments, we use the Vim-S model [73]. Moreover, we perform several ablation studies to further investigate our proposed method.
@@ 5.1 Conservation property
To verify the fulfillment of the conservation property, on which our method is based, we compare the network’s output score with the sum of relevance scores attributed to the input features, for both the GI baseline and the proposed MambaLRP. The analysis is performed for Mamba-130M models trained on the SST-2 and ImageNet datasets. Full conservation is achieved if the output score equals the sum of relevance, as indicated by the blue line in Fig. 3. Our results show that conservation is severely violated by the GI baseline, and is addressed to a large extent by MambaLRP. Residual lack of conservation is due to the presence of biases in linear and convolution layers, which are typically non-attributable.
Figure 3: Conservation property. The x-axis represents the sum of explanation scores across the input features and the y-axis shows the network’s output score. Each point corresponds to one example and its proximity to the blue identity line indicates the extent to which conservation is preserved, with closer alignment suggesting improved conservation.
@@ 5.2 Qualitative evaluation
Figure 4: Explanations generated for a sentence of the SST-2 dataset. Shades of red represent words that positively influence the model’s prediction. Conversely, shades of blue reflect negative contributions. The heatmaps of attention-based methods are constrained to non-negative values.
[…]
Figure 5: Explanations produced by different explanation methods for images of the ImageNet dataset. Explanations produced by AttnRoll and G×$\times$×AttnRoll are limited to non-negative values.
@@ 5.3 Quantitative evaluation
To quantitatively evaluate the faithfulness of explanation methods, we employ an input perturbation approach based on ranking input features by their importance [48], which can be done using either a Most Relevant First (MoRF) or Least Relevant First (LeRF) strategy. Ranked features are iteratively perturbed through a process known as flipping. We monitor the resulting changes in the output logits, $f_{c}$, for the predicted class $c$, and compute the area under the perturbation curve. The areas under the curves for LeRF and MoRF strategies are denoted by $A^{F}_{\mathrm{LeRF}}$ and $A^{F}_{\mathrm{MoRF}}$, respectively. In contrast, the insertion method starts with a fully perturbed input and progressively restores important features. The areas under the curves for this method are indicated by $A^{I}_{\mathrm{MoRF}}$ and $A^{I}_{\mathrm{LeRF}}$, for the MoRF and LeRF strategies, respectively. A reliable explanation method is characterized by low values of $A^{F}_{\mathrm{MoRF}}$ or $A^{I}_{\mathrm{LeRF}}$, and large values of $A^{F}_{\mathrm{LeRF}}$ or $A^{I}_{\mathrm{MoRF}}$. In an effort to minimize the introduction of out-of-distribution manipulations, the recent study by Blücher et al. [14] advocates for harnessing both insights to derive a more resilient metric. Therefore, we follow the same strategy as [14, 2] to evaluate explanation methods. The evaluation metrics are defined as $\Delta A^{F}=A^{F}_{\mathrm{LeRF}}-A^{F}_{\mathrm{MoRF}}$ and $\Delta A^{I}=A^{F}_{\mathrm{MoRF}}-A^{F}_{\mathrm{LeRF}}$. In both metrics, a higher score is preferable, as it signifies a more accurate and reliable explanation method.
The outcomes of this analysis are represented in Table 1. MambaLRP consistently achieves highest faithfulness scores in comparison to other baseline methods. We observe that GI struggles with noisy attributions, leading to low faithfulness scores. However, methods like Integrated Gradients and G×$\times$×AttnRoll have shown improvements in this regard. We note that LRP (LN-rule) outperforms most methods across the majority of the text classification tasks. Nevertheless, its performance is notably inferior compared to MambaLRP. Overall, we observe that MambaLRP significantly outperforms all other methods by a substantial margin. In both vision and NLP experiments, attention-based methods have shown superior performance compared to the purely gradient-based approaches. Results of the complementary insertion experiment are presented in Appendix C.4, which consistently confirm the observations from the flipping experiment.
Table 1: Evaluating explanation methods. Higher scores $\Delta{A}^{F}$ indicate more faithful explanations.
@@
Mamba 130M Mamba 1.4B Mamba 130M Mamba 1.4B Mamba 130M Mamba 1.4B Mamba 130M Mamba 1.4B Vim-S
Random -0.012 -0.106 0.044 -0.014 0.010 0.002 -0.001 0.000 0.000
GI [54] 0.078 -0.106 0.200 -0.634 -0.039 -0.039 -0.787 -0.409 0.041
SmoothGrad [55] 1.377 -0.383 1.661 -2.300 0.486 -0.687 1.808 -1.852 0.121
IG [58] 0.857 0.216 1.296 1.065 0.453 0.218 1.808 2.010 0.328
AttnRoll [4] 0.657 0.431 2.228 1.076 0.242 0.371 0.389 1.483 0.714
G $\times$ AttnRoll [4] 1.190 0.626 3.126 3.006 0.513 0.554 2.003 4.706 0.792
LRP (LN-rule, [3]) 0.877 0.961 2.217 3.456 0.673 0.656 3.079 5.199 0.647
MambaLRP (ours) 1.978 1.248 3.906 4.234 0.989 0.897 3.523 5.397 1.878
@@
Table 2: Analyzing the impact of ablating the three proposed propagation rules on $\Delta{A}^{F}$ for the components in MambaLRP.
@@
0.577 0.453
1.721 1.372
1.943 1.794
1.978 1.878
Table 3: Frequency of gendered words in explanations for ‘Nurse’ and ‘Surgeon’ classes of the Medical BIOS dataset across language models.
@@ 6 Use cases
The ability of SSMs to model long-range dependencies is considered an important improvement over previous sequence models. In this use case, we analyze the extent to which the pretrained Mamba-130M model can use information from the entire context window. We use the HotpotQA [69] subset from the LongBench dataset [11], designed to test long context understanding. After selecting all 127 instances, containing sequences up to 8192 tokens, we prompt the model to summarize the full paragraph by generating ten additional tokens. Fig. 6 shows the distribution of the positional difference between a relevant token and the currently generated token. While we observe a pronounced pattern of attributing to the last few tokens, as seen in prior language generation studies [70, 52], the extracted explanations also identified relevant tokens across the entire context window, as presented for one example in Fig. 6 (right). This suggests that the model is indeed capable of retrieving long-range dependencies. We clearly see that in order to complete the sentence and assign a year to the album release date, the model analyzes previous occurrences of chronological information and MambaLRP identifies evidence supporting the decision for the date being ‘1972’ as relevant. This demonstrates the previously speculated long-range abilities of the Mamba architecture [29].
Figure 6: Analysis of the position of tokens relevant for next token generation. Left: Distribution of absolute position of the ten most relevant tokens for the prediction of the next word. Right: Long-range dependency between tokens of the input and the predicted next token (here: 1972).
Figure 7: Explanation-based retrieval accuracy in the needle-in-a-haystack test verifying model reliance on relevant features for different context lengths.

### v3

@@
MambaLRP: Explaining Selective State Space Sequence Models
Farnoush Rezaei Jafari1,211footnotemark: 1  Grégoire Montavon3,2,1  Klaus-Robert Müller1,2,4,5,6  Oliver Eberle1,211footnotemark: 1
1Machine Learning Group, Technische Universität Berlin, 10587 Berlin, Germany
2BIFOLD – Berlin Institute for the Foundations of Learning and Data, 10587 Berlin, Germany
3Department of Mathematics and Computer Science, Freie Universität Berlin,
Arnimallee 14, 14195 Berlin, Germany
4Department of Artificial Intelligence, Korea University, Seoul 136-713, South Korea
5Max Planck Institute for Informatics, Stuhlsatzenhausweg 4, 66123 Saarbrücken, Germany
6Google DeepMind, Berlin, Germany
Correspondence to: rezaeijafari@campus.tu-berlin.de, oliver.eberle@tu-berlin.de
Аннотация
@@
38th Conference on Neural Information Processing Systems (NeurIPS 2024).
@@ 1 Introduction
Explaining the predictions of Mamba models is however challenging due to their highly non-linear and recurrent structure. A recent study [4] suggests viewing these models as attention-based models, enabling the use of attention-based explanation methods [1, 18]. Yet, the explanations produced by attention-based techniques are often unreliable and exposed to potential misalignment between input features and attention scores [75, 38]. As an alternative, Layer-wise Relevance Propagation (LRP) [10] decomposes the model function with the goal of explicitly identifying the relevance of input features by applying purposely designed propagation rules at each layer. A distinguishing feature of LRP is its adherence to a conservation axiom, which prevents the artificial amplification or suppression of feature relevance in the backward pass. LRP has been demonstrated to produce faithful explanations across various domains (e.g. [7, 62, 3, 20]). Nevertheless, the peculiarities of the Mamba architecture are not addressed by the existing LRP procedures, which may lead to the violation of the conservation property and result in unreliable explanations.
@@ 2 Related Work
Transformers [69] have emerged as the most widely used architectures for sequence modeling. However, their computational limitations, particularly with large sequence lengths, have restricted their applicability in modeling long sequences. Addressing these computational limitations, recent works [34, 35] have introduced structured state space sequence models (SSMs) as an alternative approach. SSMs are a class of sequence modeling methods, leveraging the strengths of recurrent, convolutional, and continuous-time methods, demonstrating promising performance across various domains, including language [30, 46], image [77, 13, 50], and video [71] processing, and beyond [59, 22, 42]. A recent advancement by Gu and Dao [33] introduced selective SSM, an enhanced data-dependent SSM with a selection mechanism that adjusts its parameters based on the input. Built on this dynamic selection, the Mamba architecture fuses the SSM components with multilayer perceptron (MLP) blocks. This fusion simplifies the architecture while improving its ability to handle various sequence modeling tasks, including applications in language processing [6, 53, 72], computer vision [41, 83, 78], medical imaging [44, 76, 31, 56, 40, 74, 73], and graphs [70, 14]. This fast adoption of SSMs and Mamba models underscores the need for reliable explanations of their predictions.
@@
In efforts to explain Mamba models, [51] analyzed if the interpretability tools originally designed for Transformers can also be effectively applied to architectures such as Mamba. In this context, Ali et al. [4] and Zimerman et al. [84] recently proposed viewing the internal computations of Mamba models as an attention mechanism. This approach builds upon previous works that use attention signal as explanation, including Attention Rollout [1] and variants thereof [18, 17]. While these approaches can provide some insight, they inherit the limitations of using attention as an explanation [75, 38], including their inability to capture potential misalignment between tokens and attention scores, and the limited performance in empirical faithfulness evaluations. Alternative Explainable AI methods, not yet applied to Mamba models but in principle applicable to any model, include techniques using input perturbations [81, 85, 29] or leveraging gradient information [11, 64, 68, 65, 63]. Despite their wide applicability, these models have certain drawbacks, such as requiring multiple function evaluations for a single explanation or being susceptible to gradient noise, resulting in subpar performance, as our benchmark experiment will demonstrate. Alternatively, deriving tailored approaches that reflect the underlying model structure, has shown to be a promising direction in developing better attribution methods based on gradient analysis of the prediction function [7, 27, 62, 3]. In the Layer-wise Relevance Propagation framework, this necessitates suitable propagation rules, which are currently lacking for the Mamba architecture. To tackle these challenges, we introduce MambaLRP as an efficient solution for the computation of reliable and faithful explanations that are theoretically grounded in the axiom of relevance conservation.
@@ 4.2 Relevance propagation in selective SSMs (S6)
Рис. 2: Unfolded view of SSM, highlighting two subsets of nodes, the relevance of which should be conserved throughout relevance propagation.
@@ 4.3 Relevance propagation in multiplicative gates
As for the previous examples, this ensures the conservation property $\mathcal{R}(x)=\mathcal{R}(y)$. An alternative would have been to make $y$ linear by detaching only one of the terms in the product, as done for the SiLU activation or the SSM component. However, the strategy of Eq. 9 better maintains the directionality given by the gradient. We further compare these alternatives in an ablation study presented in Appendix C.5, demonstrating empirically that our proposed approach performs better.
@@ 4.4 Additional modifications and summary
The propagation strategies developed for the Mamba-specific components complement previously proposed approaches for other layers, including propagation through RMSNorm layers [3] and convolution layers via robust LRP-$\gamma$ rules [49, 25] and their generalized variants. A summary of these additional enhancements is provided in Appendix C.2. Furthermore, our proposed propagation rules are generally applicable to other models that utilize similar components, such as multiplicative gates in recent architectures [54, 52, 45, 21].
A straightforward implementation of the propagation rules can be achieved by computing MambaLRP via Gradient $\times$ Input, where the gradient computations are modified to align with the proposed rules. The procedure consists of two main steps:
@@ 5 Experiments
To evaluate our proposed approach, we benchmark its effectiveness against various methods previously proposed in the literature for interpreting neural networks. We empirically evaluate our proposed methodology using Mamba-130M, Mamba-1.4B, and Mamba-2.8B language models [33], which are trained on diverse text datasets. The training details can be found in Appendix C.1. For the vision experiments, we use the Vim-S model [83]. Moreover, we perform several ablation studies to further investigate our proposed method.
@@
We further compare the performance of our proposed method with two attention-based approaches, Attention Rollout (AttnRoll) and MambaAttr [4], which are recently proposed for Mamba models. Both methods are extensions of techniques originally developed for Transformer models: Attention Rollout [1] and Gradient×$\times$×Attention Rollout [18].
@@ 5.1 Conservation property
To verify the fulfillment of the conservation property, on which our method is based, we compare the network’s output score with the sum of relevance scores attributed to the input features, for both the GI baseline and the proposed MambaLRP. The analysis is performed for Mamba-130M and Vim-S models trained on the SST-2 and ImageNet datasets, respectively. Full conservation is achieved if the output score equals the sum of relevance, as indicated by the blue line in Fig. 3. Our results show that conservation is severely violated by the GI baseline, and is addressed to a large extent by MambaLRP. Residual lack of conservation is due to the presence of biases in linear and convolution layers, which are typically non-attributable.
Рис. 3: Conservation property. The x-axis represents the sum of relevance scores across the input features and the y-axis shows the network’s output score. Each point corresponds to one example and its proximity to the blue identity line indicates the extent to which conservation is preserved, with closer alignment suggesting improved conservation.
@@ 5.2 Qualitative evaluation
Рис. 4: Explanations generated for a sentence of the SST-2 dataset. Shades of red represent words that positively influence the model’s prediction. Conversely, shades of blue reflect negative contributions. The heatmaps of attention-based methods are constrained to non-negative values.
Furthermore, we show the explanations produced to interpret the Vim-S model’s predictions on images of the ImageNet dataset in Fig. 5. Purely gradient-based explanations tend to identify unspecific noisy features, while both attention-based approaches, AttnRoll and MambaAttr, are more effective at highlighting significant features. Among these methods, MambaLRP stands out for its ability to generate explanations that are particularly focused on key features used by the model to make a prediction. Take, for instance, the first image classified under the ‘African elephant’ category. We can see that the explanation generated by MambaLRP not only includes all occurrences of the ‘African elephant’ object but also highlights its distinctive features, such as the tusks. In the second image labeled ‘wild boar’, despite the presence of multiple objects in the image, MambaLRP’s explanation remains focused on the ‘wild boar’ object, disregarding other objects. Moreover, in the third instance, MambaLRP uncovers a spurious correlation, the presence of a watermark in Chinese, influencing the model’s prediction, a subtlety overlooked or not fully represented by other methods. Further qualitative results can be found in Appendix C.6.
Рис. 5: Explanations produced by different explanation methods for images of the ImageNet dataset. AttnRoll and MambaAttr are limited to non-negative heatmap values.
@@ 5.3 Quantitative evaluation
To quantitatively evaluate the faithfulness of explanation methods, we employ an input perturbation approach based on ranking input features by their importance [57], which can be done using either a Most Relevant First (MoRF) or Least Relevant First (LeRF) strategy. Ranked features are iteratively perturbed through a process known as flipping. We monitor the resulting changes in the output logit, $f_{c}$, for the predicted class $c$, and compute the area under the perturbation curve. The areas under the curves for LeRF and MoRF strategies are denoted by $A^{F}_{\mathrm{LeRF}}$ and $A^{F}_{\mathrm{MoRF}}$, respectively. In contrast, the insertion method starts with a fully perturbed input and progressively restores important features. The areas under the curves for this method are indicated by $A^{I}_{\mathrm{MoRF}}$ and $A^{I}_{\mathrm{LeRF}}$, for the MoRF and LeRF strategies, respectively. A reliable explanation method is characterized by low values of $A^{F}_{\mathrm{MoRF}}$ or $A^{I}_{\mathrm{LeRF}}$, and large values of $A^{F}_{\mathrm{LeRF}}$ or $A^{I}_{\mathrm{MoRF}}$. In an effort to minimize the introduction of out-of-distribution manipulations, the recent study by Blücher et al. [15] advocates for harnessing both insights to derive a more resilient metric. Therefore, we follow the same strategy as [15, 2] to evaluate explanation methods. The evaluation metrics are defined as $\Delta A^{F}=A^{F}_{\mathrm{LeRF}}-A^{F}_{\mathrm{MoRF}}$ and $\Delta A^{I}=A^{F}_{\mathrm{MoRF}}-A^{F}_{\mathrm{LeRF}}$. For both metrics, a higher score is preferable, as it signifies a more accurate and reliable explanation method.
The outcomes of this analysis are represented in Table 1. MambaLRP consistently achieves highest faithfulness scores in comparison to other baseline methods. We observe that GI struggles with noisy attributions, leading to low faithfulness scores. However, methods like Integrated Gradients and MambaAttr have shown improvements in this regard. We note that LRP (LN-rule) outperforms most methods across the majority of the text classification tasks. Nevertheless, its performance is notably inferior compared to MambaLRP. Overall, we observe that MambaLRP significantly outperforms all other methods by a substantial margin. In both vision and NLP experiments, attention-based methods have shown superior performance compared to the purely gradient-based approaches.
Таблица 1: Evaluating explanation methods. Higher scores $\Delta{A}^{F}$ indicate more faithful explanations.
@@
Mamba 130M Mamba 1.4B Mamba 2.8B Mamba 130M Mamba 1.4B Mamba 2.8B Mamba 130M Mamba 1.4B Mamba 2.8B Mamba 130M Mamba 1.4B Mamba 2.8B Vim-S
Random -0.012 -0.106 0.007 0.044 -0.014 -0.037 0.010 0.002 $0$0.001 0.000 0.000 -0.001
GI [64] 0.078 -0.106 -0.043 0.200 -0.634 -1.434 -0.039 -0.039 0.083 -0.787 -0.409 -1.533 -0.018
SmoothGrad [65] 1.377 -0.383 -0.675 1.661 -2.300 -1.908 0.486 -0.687 -0.747 1.808 -1.852 -4.228 0.209
IG [68] 0.857 0.216 0.322 1.296 1.065 1.937 0.453 0.218 0.331 1.808 2.010 4.314 1.217
AttnRoll [4] 0.657 0.431 0.452 2.228 1.076 2.241 0.242 0.371 0.292 0.389 1.483 0.530 2.427
MambaAttr [4] 1.190 0.626 0.341 3.126 3.006 5.326 0.513 0.554 0.343 2.003 4.706 3.849 2.676
LRP (LN-rule, [3]) 0.877 0.961 0.820 2.217 3.456 5.305 0.673 0.656 0.731 3.079 5.199 5.094 2.548
MambaLRP (ours) 1.978 1.248 1.157 3.906 4.234 7.083 0.989 0.897 0.899 3.523 5.397 5.637 4.715
@@
Таблица 2: Analyzing the impact of ablating the three proposed propagation rules on $\Delta{A}^{F}$ for the components in MambaLRP.
@@
0.577 0.144
1.721 4.022
1.943 4.618
1.978 4.715
Таблица 3: Frequency of gendered words in explanations for ‘Nurse’ and ‘Surgeon’ classes of the Medical BIOS dataset across language models.
@@ 6 Use cases
The ability of SSMs to model long-range dependencies is considered an important improvement over previous sequence models. In this use case, we analyze the extent to which the pretrained Mamba-130M model can use information from the entire context window. We use the HotpotQA [79] subset from the LongBench dataset [12], designed to test long context understanding. After selecting all 127 instances, containing sequences up to 8192 tokens, we prompt the model to summarize the full paragraph by generating ten additional tokens. Fig. 6 shows the distribution of the positional difference between a relevant token and the currently generated token. While we observe a pronounced pattern of attributing to the last few tokens, as seen in prior language generation studies [80, 61], the extracted explanations also identified relevant tokens across the entire context window, as presented for one example in Fig. 6 (right). This suggests that the model is indeed capable of retrieving long-range dependencies. We clearly see that in order to complete the sentence and assign a year to the album release date, the model analyzes previous occurrences of chronological information and MambaLRP identifies evidence supporting the decision for the date being ‘1972’ as relevant. Our analysis demonstrates the previously speculated long-range abilities of the Mamba architecture [33], which we further explore in a comparison to Transformers in Appendix C.8.
Рис. 6: Analysis of the position of tokens relevant for next token generation. Left: Distribution of absolute position of the ten most relevant tokens for the prediction of the next word. Right: Long-range dependency between tokens of the input and the predicted next token (here: 1972).
Рис. 7: Explanation-based retrieval accuracy in the needle-in-a-haystack test verifying model reliance on relevant features for different context lengths.